# Re-Bottleneck: Latent Re-Structuring for Neural Audio Autoencoders

### v1

@@ 2.2 Latent Space Training
Concurrently, the Re-Bottleneck $\widetilde{\bm{z}}$ can optionally be regularized with a KL-divergence term $\mathcal{L}_{\mathrm{kl}}^{R}$ to encourage its distribution towards a desired prior, similar to the base autoencoder’s bottleneck regularization. Finally, and critically, we apply task-specific structural losses to the Re-Bottleneck $\widetilde{\bm{z}}$ These losses, denoted collectively as $\mathcal{L}_{\mathrm{task}}^{R}$, are designed to enforce the desired properties on $\widetilde{\bm{z}}$ that are relevant to the downstream task (e.g., disentanglement losses, channel-wise sparsity, equivariance penalties). The specific form of $\mathcal{L}_{\mathrm{task}}^{R}$ depends on the particular downstream objective.
@@ 2.3.3 Equivariant Re‐Bottleneck
This loss explicitly trains the filtered latent representation $h_{\alpha}(\widetilde{\bm{z}})$ to match the latent representation of the filtered input $R_{E}(A_{E}(g_{\alpha}(\bm{x})))$. Inspired by [21], we modify our reconstruction term,
@@ 3.1 Re-Bottleneck Experiments
For the base model variant, the loss weights are set as follows: $\lambda_{\mathrm{rec}}=1.0$, $\lambda_{\mathrm{kl}}=10^{-4}$, $\lambda_{\mathrm{adv}}=0.5$, and $\lambda_{\mathrm{fm}}=1$. In the semantic variant, $\lambda_{\mathrm{task}}=2.5$, and in the equivariant variant $\lambda_{\mathrm{task}}=0.5$. General hyperparameters include optimizing the encoder–decoder with Adam using a learning rate of $5\times 10^{-4}$ and the discriminator with $1\times 10^{-4}$. The InfoNCE temperature $\tau$ was set to $0.07$. In the ordered variant, the mask is only applied to 75% of the batch. We trained the ordered and equivariant variants on 3 s chunks with a batch size of 64 for 12 epochs. Due to GPU memory limitations, for the semantic variant we halved the batch size and doubled the chunk length. We train on a single NVIDIA L40S GPU for $<48$ hours.
@@ 3.4 Evaluation
We evaluate reconstruction fidelity using STFT, mel distance, and SISDR, all computed with the auraloss library [27]. In the case of the diffusion model, generation quality is measured by the FAD via the fadtk toolkit [28] configured with the CLAP-LAION-Music model. To assess prompt alignment, we report the CLAP score, which quantifies audio–text similarity [4].
@@ 4.2 Semantically Aligned Re-Bottleneck
We measure alignment using Centered Kernel Alignment (CKA) [29, 30] and Projection Weighted Canonical Correlation Analysis (PWCCA) [31] (Tab. 2). The baseline SAO bottleneck yielded 0.43 CKA / 0.63 PWCCA. A non-reconstructive upper bound (semantic loss only) achieved 0.69 CKA / 0.83 PWCCA. Our model including reconstruction loss alongside semantic objectives, achieved 0.70 CKA / 0.78 PWCCA. This shows that the model recovers nearly the upper bound’s semantic alignment, indicating strong alignment and invertibility are largely compatible in the case of BEATs. Linear probe metrics (Sec. 2.3.2) confirm that our linear probe effectively transfers semantic structure into the latent space. Reconstruction performance (Tab. 1) confirms invertibility. The semantically-aligned Re-Bottleneck showed  5% degradation versus the baseline autoencoder, trading this for a 20-60% gain in semantic capture (CKA/PWCCA). This promising trade-off is achieved with efficient training ($<48$ GPU hours).
@@
Model STFT mel SISDR
@@
Table 2: Alignment metrics (CKA / PWCCA) with mean-pooled T5 and BEATs embedding targets. “B vs T” compares the mean-pooled bottleneck representation to the target, “LP vs T” compares the linear probe to the target, and “B vs LP” compares the bottleneck to the linear probe.
@@ 4.3 Equivariant Re-Bottleneck
Table 3: (STFT / mel) distances to low-pass filtered target.
@@ 4.4 Investigating Diffusability
Fig. 4: FAD scores versus training steps, using the clap-laion-music backbone. The standard SAO (orange squares) starts best but is surpassed by the semantic Re-Bottleneck (blue circles). The ordered and T5-aligned Re-Bottlenecks do not improve downstream performance, while the equivariant matches baseline performance.

### v2

@@ 2.2 Latent Space Training
Concurrently, the Re-Bottleneck $\widetilde{\bm{z}}$ can optionally be regularized with a KL-divergence term $\mathcal{L}_{\mathrm{kl}}^{R}$ to encourage its distribution towards a desired prior, similar to the base autoencoder’s bottleneck regularization. Finally, and critically, we apply task-specific structural losses to the Re-Bottleneck $\widetilde{\bm{z}}$ . These losses, denoted collectively as $\mathcal{L}_{\mathrm{task}}^{R}$, are designed to enforce the desired properties on $\widetilde{\bm{z}}$ that are relevant to the downstream task (e.g., disentanglement losses, channel-wise sparsity, equivariance penalties). The specific form of $\mathcal{L}_{\mathrm{task}}^{R}$ depends on the particular downstream objective.
@@ 2.3.3 Equivariant Re‐Bottleneck
[…]
@@ 3.1 Re-Bottleneck Experiments
For the base model variant, the loss weights are set as follows: $\lambda_{\mathrm{rec}}=1.0$, $\lambda_{\mathrm{kl}}=10^{-4}$, $\lambda_{\mathrm{adv}}=0.5$, and $\lambda_{\mathrm{fm}}=1$. In the semantic variant, $\lambda_{\mathrm{task}}=2.5$, and in the equivariant variant $\lambda_{\mathrm{task}}=0.5$. General hyperparameters include optimizing the encoder–decoder with Adam using a learning rate of $5\times 10^{-4}$ and the discriminator with $1\times 10^{-4}$. The InfoNCE temperature $\tau$ was set to $0.07$. In the ordered variant, the mask is only applied to 75% of the batch. We trained the ordered and equivariant variants on 3 s chunks with a batch size of 64 for 12 epochs. Due to GPU memory limitations, for the semantic variant we halved the batch size and doubled the chunk length. We train on a single NVIDIA L40S GPU for <48<48 hours.
@@ 3.4 Evaluation
We evaluate reconstruction fidelity using STFT and mel-spectrogram distances and SISDR (all computed with auraloss [27]). Generation quality for diffusion models is assessed with Fréchet Audio Distance (FAD) computed via fadtk [28] configured to use CLAP-LAION-Music embeddings. Prompt alignment is measured by the CLAP audio–text similarity score [4]. Representation alignment is quantified with Centered Kernel Alignment (CKA) [29, 30] and Projection-Weighted Canonical Correlation Analysis (PWCCA) [31]. Lower is better: STFT distance, mel distance, FAD. Higher is better: SISDR, CLAP, CKA, PWCCA.
@@ 4.2 Semantically Aligned Re-Bottleneck
The baseline SAO bottleneck yielded 0.43 CKA / 0.63 PWCCA. A non-reconstructive upper bound (semantic loss only) achieved 0.69 CKA / 0.83 PWCCA. Our model including reconstruction loss alongside semantic objectives, achieved 0.70 CKA / 0.78 PWCCA. This shows that the model recovers nearly the upper bound’s semantic alignment, indicating strong alignment and invertibility are largely compatible in the case of BEATs. Linear probe metrics (Sec. 2.3.2) confirm that our linear probe effectively transfers semantic structure into the latent space. Reconstruction performance (Tab. 1) confirms invertibility. The semantically-aligned Re-Bottleneck showed  5% degradation versus the baseline autoencoder, trading this for a 20-60% gain in semantic capture (CKA/PWCCA). This promising trade-off is achieved with efficient training (<48<48 GPU hours).
@@
Model STFT ($\downarrow$) mel ($\downarrow$) SISDR ($\uparrow$)
@@
Table 2: Alignment metrics (CKA ($\uparrow$) / PWCCA ($\uparrow$)) with mean-pooled T5 and BEATs embedding targets. “B vs T” compares the mean-pooled bottleneck representation to the target, “LP vs T” compares the linear probe to the target, and “B vs LP” compares the bottleneck to the linear probe.
@@ 4.3 Equivariant Re-Bottleneck
Table 3: (STFT ($\downarrow$) / mel ($\downarrow$)) distances to low-pass filtered target.
@@ 4.4 Investigating Diffusability
Fig. 4: FAD scores ($\downarrow$) versus training steps, using the clap-laion-music backbone. The standard SAO (orange squares) starts best but is surpassed by the semantic Re-Bottleneck (blue circles). The ordered and T5-aligned Re-Bottlenecks do not improve downstream performance, while the equivariant matches baseline performance.